OpenStudy (anonymous):

hi the distance between the two foci of the hyperbola 16x^2 - 9y^2 =1 is?

OpenStudy (anonymous):

find the distance between \[\huge (\frac{ -5 }{ 12 },0),(\frac{ 5 }{ 12 },0)\]

OpenStudy (anonymous):

Hi Deena123, thanks for looking at this - could you please show me your working - I know the general equation is (x-h)^2/a^2 - (y-k)^2/b^2 =1 and the foci are at + or - ae, with e=sqrt(a^2+b^2)/a